40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 211 of 335  Not logged in ELOG logo
ID Date Author Type Category Subject
  6277   Mon Feb 13 12:02:17 2012 KojiUpdatePEMseismic BLRMS loud too

I reported the procedure to add slow channels to the FB. I guess you already have done Step.1

http://nodus.ligo.caltech.edu:8080/40m/5991

Quote:

 Tragically, this is more tricksy than I would have thought. The channels we need are "cdsEpicsOutput"s in the model.  They don't show up in Dataviewer (fast or slow channels) or the regular fast channel .ini file.  Jamie and I don't remember where these channels live and how to get them saved to frames.  I'm on top of it though.

 

  6276   Mon Feb 13 11:30:51 2012 JenneUpdatePEMseismic BLRMS loud too

Quote:

Quote:

 

 So, none of our PEM BLRMS channels are recorded as of right now.  All we have for long-term record is the StripTool on the wall.  The 0.1-0.3Hz and 0.3-1 Hz traces both show these weirdo things, but the 1Hz and up BLRMS don't have any unusual noise.

 Seems like a problem to solve on Monday so that we don't end up without trends like this again.

 Tragically, this is more tricksy than I would have thought. The channels we need are "cdsEpicsOutput"s in the model.  They don't show up in Dataviewer (fast or slow channels) or the regular fast channel .ini file.  Jamie and I don't remember where these channels live and how to get them saved to frames.  I'm on top of it though.

I did notice however, that the striptool for seismic trends is showing the wrong channels for 3-10 and 10-30 Hz.  The other 3 channels are correctly the output after the sqrt is taken, but those two (orange and red on striptool) are before the sqrt, but after the bandpass and low pass.  I'll fix that now...

  6275   Fri Feb 10 23:58:30 2012 ranaUpdatePEMseismic BLRMS loud too

Quote:

 

 So, none of our PEM BLRMS channels are recorded as of right now.  All we have for long-term record is the StripTool on the wall.  The 0.1-0.3Hz and 0.3-1 Hz traces both show these weirdo things, but the 1Hz and up BLRMS don't have any unusual noise.

 Seems like a problem to solve on Monday so that we don't end up without trends like this again.

  6274   Fri Feb 10 23:19:09 2012 kiwamuUpdateIOOcross talk causing fake seimometer signals

[ Koji / Kiwamu ]

The frequent unlock of the MC are most likely unrelated to ground motion.

Although the reason why MC became unstable is still unclear.

 

 

There are two facts which suggest that the ground motion and the MC unlock are unrelated :

(1) It turned out that the seismometer signals (C1:PEM-SEIS-STS_AAA ) have a big cross talk with the MC locking signals.

    For example, when we intentionally unlocked the MC, the seismometer simultaneously showed a step-shaped signals, which looked quite similar to what we have observed.

    I guess there could be some kind of electrical cross talk happening between some MC locking signals and the seismometer channels.

    So we should not trust the signals from the STS seismometers. This needs a further investigation.

(2) We looked at the OSEM and oplev signals of some other suspended optics, and didn't find any corresponding fluctuations.

    The suspensions we checked are ETMX, ETMY, ITMX and MC1.

     None of them showed an obvious sign of the active ground motions in the past 24 hours or so.

Quote from #6266

It seems that somehow the seismic noise became louder from about 1:00 AM.

  6273   Fri Feb 10 15:54:27 2012 steveUpdatePEMAC turned back ON

The air cond was off for 2 hrs.  I just switched it back on at 15:51

  6272   Fri Feb 10 15:52:35 2012 JenneUpdatePEMseismic BLRMS loud too

Quote:

 

 Kiwamu and Steve maybe don't know about how to trend seismic noise. If you just take the mean of the time series, you don't prove that the seismic noise got any higher. The STS has a nominally zero DC output, so the long period level shifts that you see tell you just that there was a DC offset.

This is NOT an increase in seismic noise. To see a seismic trend you should plot the trend of the BLRMS channels that we made especially for this purpose.

 So, none of our PEM BLRMS channels are recorded as of right now.  All we have for long-term record is the StripTool on the wall.  The 0.1-0.3Hz and 0.3-1 Hz traces both show these weirdo things, but the 1Hz and up BLRMS don't have any unusual noise.

  6271   Fri Feb 10 15:47:38 2012 ranaUpdatePEMseismic noise back to normal

 

 Kiwamu and Steve maybe don't know about how to trend seismic noise. If you just take the mean of the time series, you don't prove that the seismic noise got any higher. The STS has a nominally zero DC output, so the long period level shifts that you see tell you just that there was a DC offset.

This is NOT an increase in seismic noise. To see a seismic trend you should plot the trend of the BLRMS channels that we made especially for this purpose.

  6270   Fri Feb 10 15:46:59 2012 steveUpdateSUSruby wire standoff

Finally I found a company who can do Koji's improved  -hard to make-  specification on ruby or sapphire wire standoff.

NOT POLISHED excimer laser cut, wire groove radius R 0.0005" + - 0.0002"

$250 ea at 50 pieces of order

  6269   Fri Feb 10 11:46:44 2012 steveUpdateIOOseismic noise back to normal

The shaking has stopped at 9:32am  The AC was turned back on at 11:30am  We still do not have any explanation

 

Attachment 1: seism4h.png
seism4h.png
Attachment 2: seis60s.png
seis60s.png
Attachment 3: oneday.png
oneday.png
  6268   Fri Feb 10 11:01:31 2012 steveUpdateIOOcrazy ground motion

Quote:

I gave up tonight's locking activity because the MC can't stay locked.

It seems that somehow the seismic noise became louder from about 1:00 AM.

I walked around the outside of the 40-m building to see what's going on, but no one was jumping or partying.

I am leaving the MC autolocker disabled so that the laser won't be driven crazy and the WFS won't kick the MC suspensions.

 

The attachment is a 3-hour trend of the seismometer outputs and the MC trans.

MCunlock.png

 Something has started shaking last night.  Everybody is claiming to be innocent next door.

I turned off the 40m AC at 11:06

Attachment 1: seism1davg.png
seism1davg.png
  6267   Fri Feb 10 02:43:37 2012 ZachUpdateSUSOSEM LED driver noise *reduced*

I worked on the OSEM box a little more today, with the hopes of reducing the measured output current noise. I succeeded, at least modestly. It turns out that most of the noise was indeed caused by the crappy resistors.

Below is the circuit for one of the 5 LEDs. The output of the op-amp structure directly after the LT1031 reference is split between 5 stages identical to the structure on the right. I have shown just one (UR) for clarity. The various measurement points are explained below.

circuit.png

I started from the beginning of the circuit, directly after the LT1031, to make sure that the excess noise seen the other day wasn't just from a noisy reference. Below is the measured output voltage noise along with the LISO estimate. Clearly, the LT1031 is performing to spec (as it should, since it's a new part that I just put in). Note that the apparent better-than-spec performance at low frequencies is just from the AC coupling, which I needed due to the high DC level.

LT1031_vs_liso.png

Since the reference was in order, the next step was to switch out some of the crappy old resistors for nicer thin-film ones. In case anyone is interested, Frank has done some detailed investigation of excess 1/f current noise in resistors. I measured the voltage noise level at the point labeled "inter-stage measurement" above, first without any modifications and then after swapping the old 10k resistors (R1 & R2) out for nice Vishay thin-film ones. There is clearly a big improvement, and the modified circuit essentially agrees with LISO now down to 1 Hz. Below this, it looks like there could still be an issue.

interstage_vs_liso.png

I wanted to see what the improvement was in the overall output current noise of the system, so I went about measuring the current noise as I had the other day (by measuring the voltage noise across R55 and dividing by the resistance). The performance was already better than the old measurement, but not at the LISO level. So, I replaced the current-setting resistors (R54 & R55)---which were actually 3 parallel resistors on a single pad in each case---by nice Vishay ones, as well. I didn't have any that were close to the original resistance of ~287 ohms, so I put three 1k ones in parallel. This of course shifts the resistance up to 333 ohms, but that only causes a ~16% change in current. I was sure to convert voltage noise into current noise with this new resistance, though.

With this change, the total output current noise is now very close to the LISO estimate as well down to ~1 Hz.

LED_driver_vs_liso.png

Some notes:

  • First, I apologize for the noise margin at higher frequencies. I redid the higher frequency measurements with an SR560 as a preamp, but I must have screwed up the calibration because the data don't match up quite right with the LF measurements. It was clear while I was taking them that they followed the LISO trace.
  • There still seems some excess noise below 1 Hz. It could be that the noisy resistors in the parallel stages were somehow still contaminating the cleaned-up channel. I'll look into this more soon.
  6266   Fri Feb 10 02:35:29 2012 kiwamuUpdateIOOcrazy ground motion

I gave up tonight's locking activity because the MC can't stay locked.

It seems that somehow the seismic noise became louder from about 1:00 AM.

I walked around the outside of the 40-m building to see what's going on, but no one was jumping or partying.

I am leaving the MC autolocker disabled so that the laser won't be driven crazy and the WFS won't kick the MC suspensions.

 

The attachment is a 3-hour trend of the seismometer outputs and the MC trans.

MCunlock.png

  6265   Thu Feb 9 20:01:02 2012 ranaSummaryElectronicsUsing RF LP filters as dispersion units for the MFD

 WE currently use long cables to give us the dispersion that we want for the MFD. A cable gives a long delay - both the phase delay and the group delay.

But we only need the dispersion (group delay). We can get this by just using a very sharp low pass filter and having the corner be above the frequency that we have the beat signal.

For example, the MiniCircuits SLP-200+ has got a corner frequency of 200 MHz and a group delay of ~10 ns (like a 3 m vacuum delay). So we would have to use 10 of these to get the delay we now get. The passband attenuation is only 0.5 dB, so we would lose 5 dB. The cost is $35 ea. We have a few on the shelf.

OTOH, if we tune the beat frequency down to 30 MHz, we can use the SLP-30 which has a group delay of 30 ns around 30 MHz. That's like 9m at light speed. We could easily get a nice result by just using 4 or 5 SLP in series.

So why is Kiwamu using cables?? And how should we really choose the beat note frequency??

  6264   Thu Feb 9 17:11:05 2012 steveUpdateSUSmore OSEM problems

These observations of the OSEMs  were taken while taking transfer functions of oplev YAW at excitation amplitude 0.1

 

 Atm1,  C1:SUS-ETMX_SENSOR_SIDE cross coupling

Atm2,   C1:SUS-ITMX_SENSOR_LL   not excitable

Atm3-4,   BS and PRM  insensitive

 

Good OSEM list: ITMY, ETMY and SRM

 

Attachment 1: ETMXosemEXC0.1ampl.png
ETMXosemEXC0.1ampl.png
Attachment 2: IMTX-YAW0.1.png
IMTX-YAW0.1.png
Attachment 3: BSosemEXC.1ampl.png
BSosemEXC.1ampl.png
Attachment 4: PRMosemEXC.1ampl.png
PRMosemEXC.1ampl.png
  6263   Thu Feb 9 16:46:02 2012 steveUpdateSUS ALL oplev YAW transfer functions

Quote:

SUS- BS, ITMX, ITMY, PRM, SRM, ETMX & ETMY_OLPIT transfer funtion with sine wave excitation 0.1 amplitude:

 

OL_YAW transfer functions are here.

 I had two PHDs helping me to overlap the EXML files in DTT. We failed. This job requires professorial help.

Attachment 1: BS-OLYAW.1ampl.png
BS-OLYAW.1ampl.png
Attachment 2: ITMX-OLYAW.1ampl.png
ITMX-OLYAW.1ampl.png
Attachment 3: ITMY-OLYAW.1ampl.png
ITMY-OLYAW.1ampl.png
Attachment 4: PRM_OLYAW.1.png
PRM_OLYAW.1.png
Attachment 5: SRM_OLYAW.1ampl.png
SRM_OLYAW.1ampl.png
Attachment 6: ETMX-OLYAW.1ampl.png
ETMX-OLYAW.1ampl.png
Attachment 7: ETMY-OLYAW.1ampl.png
ETMY-OLYAW.1ampl.png
  6262   Thu Feb 9 13:04:11 2012 KojiSummaryIOOPMC/IMC alignment

There has been no lock of input MC for more than 5days. WTF???

I have fixed a loose mirror of the PMC input and the alignment of the MC2 Yaw.


- The PSL mech-shutter was closed. It has been opened.

- Then, I checked the MC suspensions. Mainly MC2 Yaw has kept drifting. (Fig.1)
In fact, there was no WFS actions during this drift.

Anyway, now MC2 Yaw was aligned and the lock was restored.

- It was very unsatisfactory for me that the PMC alignment kept drifting.
The trend of the PMC REFL and PMC TRANS for a year suggests that:

  • Occasional drifts exist since a year ago (or probably since we built this new PMC setup in 2010 Sep)
  • The drift has got frequent for the recent four months, and got more for the recent one month.

I went into the PMC setup and tapped several optics in order to find any loose optic.
Immediately I found that the mirror before the AOM was loose. Basically any mild tapping was enough
to misalign the mirror such that the caivty loose the TEM00 mode.

I tightened the retainer set screw of the optic and aligned the PMC again. It looks OK now as I can not
misalign this optic by the tapping anymore. But if it still remains drifting, we need to replace the mount.

Attachment 1: Untitled.png
Untitled.png
Attachment 2: Untitled2.png
Untitled2.png
  6261   Thu Feb 9 12:38:15 2012 ZachUpdateSUSOSEM LED driver noise

Frank pointed out to me that I had dumbly forgotten to include the voltage reference's noise. The LT1031 has an output noise level of ~125 nV/rHz above 10 Hz or so, and this at least makes the estimate much closer. I had also not included an extra LT1125 stage between the reference and the other stages. I guess I was tired.

LED_driver_vs_liso.png

The estimate is now within a factor of a few of the measured level, and it has roughly the right shape. Around 1 Hz, it looks like the measured data begin to roll up away from the model, though it's tough to say due to the effect of the AC coupling on the analyzer less than a decade below. If there is indeed extra noise here, Frank thinks it could be due to resistor current noise.

I'll switch one or two out for nicer ones and see if things change.

  6260   Wed Feb 8 16:37:02 2012 steveUpdateSUSITMX OSEM LL is sick

I'm driving C1:SUS-ITMX_OLYAW and PIT_EXC with amplitude 0,1-0.3 while taking transfer funtions of oplev.

The transfer functions are normal. However I noticed that the LL osem is not responding to this excitations

Healthy sensor respons should be like Atm3

Attachment 1: ITMXsensorLL.png
ITMXsensorLL.png
Attachment 2: ITMXsensorLLyaw.1.png
ITMXsensorLLyaw.1.png
Attachment 3: ETMX.2amplPIT2.png
ETMX.2amplPIT2.png
  6259   Tue Feb 7 16:30:46 2012 steveUpdateSUS ALL oplev PIT transfer functions

SUS- BS, ITMX, ITMY, PRM, SRM, ETMX & ETMY_OLPIT transfer funtion with sine wave excitation 0.1 amplitude:

 

Attachment 1: BS_PIT.png
BS_PIT.png
Attachment 2: ITMX_PIT.png
ITMX_PIT.png
Attachment 3: itmy_pit.png
itmy_pit.png
Attachment 4: PRM_PIT.png
PRM_PIT.png
Attachment 5: SRM_PIT.png
SRM_PIT.png
Attachment 6: ETMX_PIT.png
ETMX_PIT.png
Attachment 7: ETMY_PIT2.png
ETMY_PIT2.png
  6258   Tue Feb 7 03:05:08 2012 ZachUpdateSUSOSEM sat amp measurements

I did some more investigation on the OSEM box today.

Troubleshooting:

After removing some capacitors and still finding that the +15V rail was at over +20V, I decided to see if the TO-3 7815 that I removed behaved properly all by itself. It did. After some more poking around, I discovered that whoever assembled the board isolated the case of the regulator from the board. It is through the case that this package gets its grounding, so I removed the mica insulator, remounted the regulator, and all worked fine.

Since I had gotten a spare from Downs, I also replaced the LT1031 (precision 10-V reference), for fear that it had been damaged by the floating voltage regulator.

 

Noise measurements:

sat_amp_measurement_setup.png

LED Driver

With the above out of the way, I was finally able to take some measurements. The first thing I did was to look at the LED drivers. I fixed the one stage that I mentioned in my last post by adding two 820-ohm resistors in parallel with the 1k, such that it was very close to all the others (which are 806 || 806 || 1k). With that, using a red LED, I measured a current of 34.5 mA (+/- 0.1) out of each of the 5 stages (UL, UR, LL, LR, S).

I then measured the current noise of each one by monitoring the voltage across the 287-ohm resistor in series with the LED. The driver works by putting the LED in the feedback path of an inverting amp. There is a 10-V input from the LT1031, and the values of the input and feedback resistors determine the current drawn through the LED. There is a buffer (LM6321) in the path to provide the necessary current.

The LISO model I made according to that description seems to make sense. I simply modeled the LED as a small resistor and asked LISO for the current through it. The transfer function shows the proper DC response of -49.15 dB(A/V) -->  34.8 mA @ 10 V, but, the estimated current noise doesn't add up with the measured levels:

LED_driver_vs_liso.png

I have to get to the bottom of this. Two possibilities are: 1) The buffer adds noise, and/or 2) I am modeling this invalidly.

PD Amp

I also began measuring the PD amplifier noise levels, though I only measured two of them for lack of time. I find it odd that there is a 100-ohm input series resistor on what I thought would be just a transimpedance amplifier. For that reason, I want to look into how the OSEMs are connected to this guy.

In any case, I measured the output noise of two of the PD amps by shorting the input side of the 100-ohm resistors to ground, and then I divided by their TF to get the input noise level. Here it is compared with the LISO estimate. I have plotted them in units of voltage noise at the input side of the resistors for lack of a way to infer the equivalent photocurrent noise level.

pd_amp_vs_liso.png

Above 2 Hz or so, the measured level agrees with the prediction. Below this, the measured noise level increases as 1/f, while it should go as the standard 1/sqrt(f) (the manufacturer-quoted 1/f corner is at 2 Hz). Another thing to get to the bottom of.

  6257   Mon Feb 6 17:06:11 2012 steveUpdateSUSoplev transferfunction of PIT: ETMY & ETMX

Kiwamu showed me how to do transferfunction of oplev pitch

Attachment 1: ETMX_PIT.png
ETMX_PIT.png
Attachment 2: ETMY_PIT.png
ETMY_PIT.png
  6256   Mon Feb 6 11:07:21 2012 steveUpdateSUSOplev QPD Sum Trends are suspicious

Quote:

The attached trend shows a problem with the QPD sums.

Why are ETMX and ITMX so much lower than ETMY and ITMY? Are the laser's dying? Or is it the gain inside the QPD? Or the reflectivity of the coatings?

Steve - please check on Monday the laser powers and the ETM/ITM reflectivity for HeNe lasers. Maybe we have to increase the transimpedance gain in the heads.

ETMX and ETMY have 0.2 mW returning to their QPDs........so the gain must lower at  ETMX

ITMX laser 1103P has only 0.67 mW output and 0.025 mW returning to the QPD.

ITMX and ETMX oplev lanching paths  have lenses without AR coating. This is my fault. I will buy them.

          ETMX   0.2 mW           900 counts

         ITMX     0.025           1300

         SRM      0.04            2600

          BS      0.05            3500

          PRM     0.06            4000

         ETMY     0.2             9000

        ITMY      0.3            14500

Attachment 1: oplevsums.png
oplevsums.png
  6255   Fri Feb 3 23:19:22 2012 ZachUpdateSUSOSEM testing begins

I took one of the spare OSEM satellite amps (schematic) from the cabinet down the Y arm this afternoon to begin testing. I spent most of the day amassing the melange of adapters and connectors I needed to talk to the relic. The most elusive was the über-rare 64-pin IDE connector, for which neither the 40m nor Downs or Bridge had a breakout (despite there being several Phoenix boxes on each electronics rack at the 40m---hmm...). The solution I came up with was to make a breakout cable myself, only there was no 64-pin ribbon. So, I carefully fed a 50-pin and most of a 16-pin ribbon side by side into one push-down connector, and that was that:

IDE64_homestyle.png

I also finally found a 25-pin D-sub breakout just after figuring out the proper pinout for a 25-to-9 adapter, which I thought I was going to have to use. OH WELL.

Science

The first thing I figured I'd do is measure the LED drivers' current noise and see how it compared with LISO. I powered the box up and found that the TO-3 7815 regulator was putting out +20V---bad. I assumed it was broken, so I got another one from Downs and replaced it. Powered it up again and the output was still at +20V (WTF?). My suspicion is that one of the shielding capacitors has failed in some bizarre way, but I didn't have time to check this before I was beckoned to another task. This is where I'll start again next.

Another thing Frank and I noticed as we were figuring out how the driver worked was that the current-specifying resistor of one of the driver stages had not been properly modified along with the others, so it was forcing the feedback loop to rail. This mod was done precariously by adding two perpendicular sandwiched "Radd" resistors on top of the main one, so it's also possible that the ones for this stage had just been knocked off somehow (perhaps by the massive gender-switching ribbon chain hanging down on it). Steve and I noticed that there was a label on the box complaining that some part of the amp for one of the OSEMs wasn't working, but we peeled it off and threw it away because he figured it was outdated.

Anyway, in short, the plan going forward is as follows:

  • For this box
    • Measure the LED driver and PD transZ amp noises with dummy components
    • Compare with LISO to make sure they make sense
    • Measure the noise again with an OSEM connected
    • Based on the above and more LISO modeling, decide if it's a good idea to replace the LT1125's with OP497's
    • Increase the dynamic range by allowing +/-10V output, rather than +/-2V as was needed for old ADCs
  • After
    • Systematically mirror the changes in all other boxes by switching one out at a time

Comments welcome.

  6254   Fri Feb 3 20:58:56 2012 SteveConfigurationGeneralIlluminator Picture
Attachment 1: P1080526.JPG
P1080526.JPG
  6253   Fri Feb 3 20:19:38 2012 ranaUpdateSUSOplev QPD Sum Trends are suspicious

The attached trend shows a problem with the QPD sums.

Why are ETMX and ITMX so much lower than ETMY and ITMY? Are the laser's dying? Or is it the gain inside the QPD? Or the reflectivity of the coatings?

Steve - please check on Monday the laser powers and the ETM/ITM reflectivity for HeNe lasers. Maybe we have to increase the transimpedance gain in the heads.

Attachment 1: a.png
a.png
  6252   Fri Feb 3 19:57:26 2012 ranaConfigurationSUSPSL Shutter closed for SUS hysteresis test

Fri Feb 03 19:57:20 2012

Fri Feb 03 20:25:19 2012 :   Aligned all SUS to center their OL beams

Fri Feb 03 20:29:21 2012:    Aligned all SUS to make OL_PIT = 0.5

  6251   Fri Feb 3 19:54:21 2012 ranaConfigurationIOOMC trans awry

As usual, I noticed several bad things within 30 seconds of sitting in front of the workstation. Today its that there are OFF or missing filters on the MC TRANS.

is this the normal state? Screenshot attached.

Attachment 1: mctrans.png
mctrans.png
  6250   Fri Feb 3 17:31:09 2012 steveUpdateGeneralgreen welding glass

Schott, green welding glass, shade 14, 3 mm thick  was measured in the beam path of 1.2W, S polarization of 1064nm at ~1 mm diameter size as MC reflected path.

Absorption 95%, R 5% at incident angle 25-50 degrees. It looks like the perfect material for beam trap.

 

Attachment 1: 02031202.JPG
02031202.JPG
  6249   Fri Feb 3 17:29:28 2012 DenUpdateComputersc1lsc kernel

The reason I've killed the c1lsc kernel was the following - when the code starts to run, it initializes some parameters and this takes ~0.2 msec per dof. Now, the old code did nothing with a DOF if C1:OAF-ADAPT_???_ONOFF == OFF. My code still initialized the parameters but then does nothing because no witness channels are given. But it spends 8*0.2 = 1.6 msec for initializing all 8 dof. As the code is called with frequency 2k, this was the reason for crashing. Now I've corrected my code, it compiles, runs and does not kill c1lsc. However, the old code would also kill the kernel if all DOF are filtered. So, when we'll use all 8 DOF, we'll have to split variable initialization.

But this is not the biggest problem. C1OAF model must be corrected, because, as for now, all 8 DOF call the same ADAPT_XFCODE function. As this function uses static variables, they will be all messed up by different DOF signals.

  6248   Fri Feb 3 17:17:47 2012 DenUpdateIOOMC SUS misalignment

Quote:

Reminder / Moral:  Everything cannot be considered to be "working fine" if the MC isn't locking.  See if you can figure out why, and especially if it's something that you screwed up, either fix it, or better yet, ask for help and learn how to fix what you broke.

When I left this morning, Steve was still working with the MC and it was unlocked anyway, I could not check it. By "fine" I meant only watchdogs. The thing is that before starting to work with c1lsc I turned off all the coils. Crazyness that Steve saw was after I turned them on back after reboot. This is a confusing thing - restarting models on c1lsc and burt restoring them is not enough. After I did it, everything at the STATUS MEDM screen was green, but the C1:SUS-???_??PD_VAR values went up after turning on coils. So sus and lsc communicated in a bad manner after the reboot. After restarting x02 model, the watchdogs were fine again.

  6247   Fri Feb 3 16:13:49 2012 steveUpdatePEMLED lights for chamber illumination

Cold LED lights replaced hot halogen ones. Flat LED MYAL 6S,  model #112560002  24VAC

This is a LATE ENTRY.  They were purchased  in Jan 2010 and installed 6 of them around May 2010

Attachment 1: P1080526.JPG
P1080526.JPG
  6246   Fri Feb 3 15:49:10 2012 JenneUpdateIOOMC SUS misalignment

[Jenne, Den]

We moved the MC approximately back to where the sensors for each optic used to be (mostly touching MC2, but a little bit of MC1 to help the refl get back to its max value).  MC is now locked, and with the help of the WFS it's back to nominal.  I forgot to disable the WFS, so I think we aren't perfectly aligned, but we're close enough for the WFS to get us the rest of the way.  We're heading over to JClub right now, so we're going to leave it as-is.

  6245   Fri Feb 3 14:47:51 2012 steveUpdatePEMlaser interlock drawing

 Rough draft of      updated interlock drawing by Ben is here.

 

  6244   Fri Feb 3 14:44:33 2012 JenneUpdateIOOMC SUS misalignment

The mode cleaner is a little misaligned in pitch, and is very misaligned in yaw.  The lowest order mode that is flashing is TEM11.

I had a look-see at the SUS sensors, to see if there were any big jumps.  There were moderately sized jumps on all 3 mode cleaner optics.

SUSjump_3Feb2012.png

The MC's lockloss was at ~8:22am this morning, and went along with a giganto kick to the optics.  Steve tells me that Den might have been kicking up optics while doing computer things this morning, before Steve reminded him to shut off the watchdogs.  However, Steve was also taking phots/measuring things near MC Refl, so maybe he's not totally absolved of blame.  But this really looks like the optics settled to different places after big kicks.

I'm going to try to align the MC mirrors to get back to the sensor numbers from early this morning before chaos began.

Reminder / Moral:  Everything cannot be considered to be "working fine" if the MC isn't locking.  See if you can figure out why, and especially if it's something that you screwed up, either fix it, or better yet, ask for help and learn how to fix what you broke.

  6243   Fri Feb 3 10:48:24 2012 DenUpdateComputersc1lsc kernel

This morning I killed again c1lsc kernel with the new realization of fxlms algorithm. It works fine with gcc compiler during the tests. However, smth forbidden for the kernel is going on. I'll spend some more time on investigatin it. Interesting thing is that I did not even pressed "On" at the OAF MEDM screen to make the code running. c1lsc suspended even before. May be there is some function-name mismatch.

After c1lsc suspention I recomiled back non-working code and rebooted c1lsc. c1sus is also bad after c1lsc reboot as they communicate. I killed x04, lsc, ass, oaf models on the c1lsc computer and sus, mcs, rfm, pem on the c1sus computer. Then I restarted x02 model and restored its burt snapshot from 08:07. After I started all models back and restored their burt snapshots from 08:07. Then I diag reset all started models.

Before starting new fxlms code I've shutted down all the optics so that possible c1lsc suspention would not make them crazy. After reboot I turned the coils back. Everything seems to work fine.

  6242   Wed Feb 1 17:00:57 2012 steveUpdateIOOlaser is back ON

Quote:

 

The 2W PSL laser is turned off.  The danger laser lights are not illuminated at the entry doors because of malfunctioning electronic circuit!!!

Laser safety glasses are still required!  Other lasers are in operation!

 BEN fixed the interlock.  The laser is turned ON. Thanks for all, Rich and Sam who came over to help. Atm1

All emergency shut- off switches, lights and door indicators are working at this moment. More about this tomorrow.

Atm2, PSL enclosure interlock jungle without REAL schematic drawing.....at this point.... We all agreed it is easier to redo the hole thing than find the problem

Atm3, Emergency shut off switches and illuminated signs from  entry doors to AC on-off box  ( Use this switches in emergency ONLY,  otherwise leave alone , even it is labeled obsolete !)

Summery: I still do not really know what was wrong.

 

Attachment 1: P1080525.JPG
P1080525.JPG
Attachment 2: P1080514.JPG
P1080514.JPG
Attachment 3: P1080518.JPG
P1080518.JPG
  6241   Tue Jan 31 17:13:49 2012 steveUpdateIOOonly the PSL laser is off

 

The 2W PSL laser is turned off.  The danger laser lights are not illuminated at the entry doors because of malfunctioning electronic circuit!!!

Laser safety glasses are still required!  Other lasers are in operation!

  6240   Tue Jan 31 14:58:30 2012 kiwamuUpdateIOOlaser shuts down

[Steve/ Kiwamu]

 We found that the laser had completely shut off for ~ 4 hours even with all the PSL doors closed.

We are guessing it is related to the interlock system and Steve is working on it to fix it.

Quote from #6239

 The 2W Innilight shutdown shut when I opened side door for safety scan. This was not a repeatable by opening -closing side doors later on. Turned laser on, locked PMC and MC locked instantly. The MC was not locked this moring and it seemed that the MC2 spot was still some high order mode

like yesterday. MC lock was lost when the janitor bumped something around the MC.

 

  6239   Tue Jan 31 08:44:10 2012 steveUpdateIOOlaser shuts down

 

 The 2W Innilight shutdown shut when I opened side door for safety scan. This was not a repeatable by opening -closing side doors later on. Turned laser on, locked PMC and MC locked instantly. The MC was not locked this moring and it seemed that the MC2 spot was still some high order mode

like yesterday. MC lock was lost when the janitor bumped something around the MC.

Attachment 1: laseroff.png
laseroff.png
  6238   Mon Jan 30 23:10:02 2012 kiwamuUpdateSUSdrift mon script needs to be fixed

[Rana / Kiwamu]

We tried to set some parameters for the suspension drift monitor but the old matlab script, which automatically sets the values, didn't run because it uses the old mDV protocol.

The attached link below is a description about the script.

https://wiki-40m.ligo.caltech.edu/Computers_and_Scripts/All_Scripts#Drift

It needs to be fixed or upgraded by pynds.

  6237   Mon Jan 30 16:18:51 2012 steveUpdatePEMRoscolux colored film transmittance at 1064 nm

 

 Roscolux filter films  #74 night blue,  0.003" thick  and  #26 light red, 0.002" thick  were measured in the beam path of  ~6 mm diameter,  1W 1064 nm .

T 90%  + - 5% at 0-30 degrees of  incident angles and R ~10 % 

These sandwitched thin films of policarbonate-polyester filters are not available in thicker forms. Rosco is recommending them to be cooled by air if used in high power beam.

These filters did not get warm at all in 1W, so absorption must be very small.

  6236   Mon Jan 30 08:17:06 2012 steveUpdateIOOPMC

Quote:

I have realigned the beam pointing to PMC. The transmitted light increased from 0.74 to 0.83.

The misalignment was mainly in pitch.

 The PMC pointing has changed, so MC is resonating in high order modes.

Attachment 1: sickPMC.png
sickPMC.png
  6235   Fri Jan 27 17:16:05 2012 kiwamuUpdateLSChypothetical glitch scenario

Here is a hypothetical scenario which could make the glitches in the LSC error signals. It can be considered as a 4 step phenomenon.

        (1) up conversion noise due to a large motion at 3 Hz

   => (2) rms level exceeds the line width (a.k.a. linear range) in some LSC sensors

   => (3) unlocks some of the DOFs  in a moment

   => (4) glitches due to the short unlock.

- - plan  - -

 In order to check this hypothesis the low finesse PRMI must serve as a good test configuration.

What I will do is to gradually decrease the offset in MICH such that the finesse of PRMI becomes higher.

And at each different finesse I will check the spectra, glitch rate, and etc.

Quote from #6231

        Low finesse PRMI      

In this configuration NO glitches ( a high speed signal with an amplitude of more than 4 or 5 sigma) were found when it was locked.

Is it because I didn't use AS55 ?? or because the finesse is low ??

Also, as we have already known, the up conversion noise (#6212) showed up -- the level of the high frequency noise are sensitive to the 3 Hz motion.

  6234   Fri Jan 27 16:55:28 2012 JenneUpdateGreen LockingY-green realigned

The Yarm green laser really wanted to lock on a 01/10 mode, so Kiwamu suggested I go inside and realign the green beam to the arm.  I did so, and now it's much happier locked on 00 (the Yarm is resonating both green and IR right now).

  6233   Fri Jan 27 13:13:03 2012 JenneUpdateSUSITMs tripped

Sitting down to start cavity measurements, I found both ITMs tripped.  It must have happened a while ago (I didn't bother to check dataviewer trends) because both had rms levels of <5 counts, so they've had a while to sit and quiet down.

  6232   Fri Jan 27 09:05:41 2012 steveUpdateSUSOSEM locking plunger

Quote:

Our existing 300 series SS plungers from McMastercar #8476A43 are silver plated as Atm2 shows.

Problems:  1, they become magnetized after years being close to the magnets

                     2, they oxidize by time so it is hard to turn them

                    

I looked around to replace them.

Titanium body, nose and beryllium copper spring. None magnetic for UHV enviorment.

Can be made in 7 weeks at an UNREASONABLE $169.00 ea at quantity of 50

 In order to get a better price from Vlier's Tom Chen I changed Ti body back to SS304L-siver plated and music wire spring. The price is still ~$120 ea. at quantity 50

I will talk to Mike G about modifying the  McMaster plunger with a hex nut.

  6231   Fri Jan 27 06:07:47 2012 kiwamuUpdateLSCglitch hunting

I went through various IFO configurations to see if there are glitches or not.

Here is a summary table of the glitch investigation tonight. Some of the cells in the table are still not yet checked and they are just left blank.

IFO configuration  Yarm
Xarm
MICH
Half PRMI
low finesse PRMI
PRMI (carrier)
PRMI (sideband)
DRMI
AS55 NO NO NO   up conversion noise glitch glitch glitch
REFL11 NO NO NO   up conversion noise
glitch glitch glitch
REFL33 NO NO NO   - glitch glitch glitch
REFL55 NO NO NO   up conversion noise
glitch glitch glitch
REFL165 NO NO NO   - glitch glitch glitch
POX11 - NO NO     glitch glitch glitch
POY11 NO - NO     glitch glitch glitch
POP55 - -            
                 

 

        Low finesse PRMI      

The low finesse PRMI configuration is a power-recycled MIchelson with an intentional offset in MICH to let some of the cavity power go through MICH to the dark port.

To lock this configuration I used ASDC plus an offset for MICH and REFL33 for PRCL.

The MICH offset was chosen so that the ASDC power becomes the half of the maximum.

In this configuration NO glitches ( a high speed signal with an amplitude of more than 4 or 5 sigma) were found when it was locked.

Is it because I didn't use AS55 ?? or because the finesse is low ??

Also, as we have already known, the up conversion noise (#6212) showed up -- the level of the high frequency noise are sensitive to the 3 Hz motion.

  6230   Fri Jan 27 05:21:43 2012 kiwamuUpdateGreen Lockingfine alignment of the Y end green setup

I did a fine alignment on the Y end green setup. The green light became able to be locked again.

Quote from #6227

 The alignment is finished after the realization that the 3rd steering mirror had to be adjusted too.

  6229   Thu Jan 26 19:28:02 2012 kiwamuUpdateIOOPMC low transmission

After I recovered the lock of PMC, I found that the PMC transmission was quite low. It was about 0.26 in the EPICS display.

I zeroed the PSL temperature feedback value which had been -2.3 and then the PMC transmission went back to a normal value of 0.83.

I believe it was because the PSL was running with two different oscillation modes due to the big temperature offset.

  6228   Thu Jan 26 15:35:23 2012 JenneUpdateIOOMC ~1Hz badness

The mode cleaner is super unhappy.  It's rocking around at ~1Hz.

I turned off the WFS and turned them back on after the MC was locked, and it seems a little happier now.  At least it's not falling out of lock ~1/minute.

ELOG V3.1.3-